OpenStudy (anonymous):

someone help please last quetion been at it for5 hours fan and medal A quadratic equation is shown below: 3x2 − 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (5 points) Part B: Solve 9x2 + 3x − 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (5 points)

OpenStudy (anonymous):

@ganeshie8 can u pleease help me with this last 1

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

For part b, factorization will work. 9x^2 + 3x - 2 = 0 9x^2 + 6x - 3x - 2 = 0 3x(3x + 2) - 1(3x+2) = 0 (3x-1)(3x+2) = 0 x = 1/3, -2/3

OpenStudy (anonymous):

For A use this: Quadratic equation: \(ax^2+bx+c=0\) Discriminant of quadratic formula: \(b^2-4ac\) a = 3, b = -16, c = 2

OpenStudy (anonymous):

If \(b^2-4ac\) is positive → 2 real solutions 0 → 1 real solution negative → 0 real solutions

OpenStudy (anonymous):

thank you guys so much

OpenStudy (anonymous):

@LegendaryNikki thanks

OpenStudy (anonymous):

yw